# Imbalanced Data Clustering using Equilibrium K-Means

### v2

@@ V Numerical Experiments
The maximum number of iterations is set to 100. FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [33]. The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothing parameter of EKM is set to $\alpha=1$. We evaluate EKM’s performance on 12 datasets, including two synthetic and ten real datasets. Besides, we also apply EKM to the application of image segmentation. Because of space limitations, experiment results from six datasets and the image segmentation are shown in the paper. Full experiment results can be found in the supplemental material. The six datasets and the test image are introduced below.
V-A Datasets
V-A1 Synthetic 2-D Ball
We generate data from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$, and the same identity covariance matrix. The dataset consists of 2,100 samples, of which 2000 samples from the normal distribution with means $(-2,2)$, and 100 samples from the other two normal distributions (each distribution for 50 samples). The clustering results are shown in Fig. 4.
V-A2 Synthetic 2-D Noisy Ball
We add 100 uniformly distributed noise data (around 5% of the number of clean samples) to the synthetic 2-D ball dataset in order to test the algorithms’ robustness against noise. The clustering outcomes are displayed in Fig. 5.
V-A3 Fisher’s Iris Data
The Fisher’s Iris dataset [34, 35] is a well-known dataset for evaluating the performance of clustering algorithms. It comprises 50 samples from each of three iris species (Iris setosa, Iris virginica, and Iris versicolor). Two features, the width and the length of the sepals, are utilized as clustering features and normalized to have zero means and unit variance. Fig. 6 depicts the clustering results.
V-A4 Imbalanced Fisher’s Iris Data
Because the instance distribution in the Iris dataset is balanced. To examine the capability of imbalanced data clustering, we generate an imbalanced Fisher’s Iris dataset by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class. The clustering results of this imbalanced dataset are presented in Fig. 7.
V-A5 Wisconsin Diagnostic Breast Cancer
Wisconsin Diagnostic Breast Cancer (WDBC) [36] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses. This database contains 569 instances, with an instance distribution of 357 benign and 212 malignant cases. We use the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance. Fig. 8 presents the clustering results where the mean radius and the mean texture are used as coordinates for visualization.
V-A6 Imbalanced Wisconsin Diagnostic Breast Cancer
The data distribution in WDBC is quite balanced. To examine the ability of imbalanced data clustering, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a distribution of 357 benign and 12 malignant instances. We still use the first three features for clustering, each normalized with zero mean and unit variance. The clustering results are shown in Fig. 9.
V-A7 Color-based image segmentation
The test image coloredChips.png is chosen from Matlab’s image gallery (shown in Fig. 10a). The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface). Clustering is conducted in the RGB color space and the segmentation results can be viewed in Fig. 10b - 10e.
V-B Discussion of Results
In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data. In particular, HKM, FKM, and MEFC suffer from the uniform effect when data has an imbalanced distribution. The csiFKM and siibFKM algorithms are sensitive to noise data: Fig. 5c and Fig. 5d show that their centroids are highly deviated from the correct positions in the presence of noise. In comparison, EKM is more robust to noise because of the anchoring effect produced by the positive gradients (see Section IV-B), as seen in Fig. 5e.
The strong performance of EKM shown in Fig. 9 demonstrates its ability to identify clustered anomalies. This makes up a missing function of HKM and FKM in anomaly detection: they are typically used to detect isolated anomalies, not clustered anomalies [37, 38, 39, 40].
The results of color-based image segmentation are shown in Fig. 10b - 10e. HKM, FKM, and MEFC erroneously separate the background (the desk surface) into two clusters. This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within each clusters. On the other hand, EKM produces the expected segmentation result.
Figure 4: (a) Synthetic 2-D ball data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
Figure 5: (a) Synthetic 2-D noisy ball data. (b) Clustering by FKM. (c) Clustering by csiFKM. (d) Clustering by siibFKM. (e) Clustering by EKM.
Figure 6: (a) Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
Figure 7: (a) Imbalanced Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
[…]

### v3

@@ V-A Experimental Setup
We set the maximum number of iterations to 500 to prevent algorithms from infinitely iterating due to failure to converge. Except for sporadic cases, all algorithms can reach the specified convergence within 500 iterations.
Centroids are initialized using the K-means++ algorithm [39]. Given that each run of K-means++ yields different outputs and the objectives of the tested clustering algorithms are known to be non-convex (i.e., multiple local optima exist), convergence may happen at one of the local optimal points. A common way of finding the global optimum is to carry out a number of replications followed by a selection of the best (lowest) objective value. Hence, each trial includes 100 repetitions, and we select the repetition with the lowest objective value as the final result for that trial. We average the performance of 50 trials to ensure the rationality of the experiment. This means we conduct $50\times 100=5000$ runs for each algorithm. 100 repetitions in each trial is still within the practical executable range and each trial can give almost consistent results as indicated by a small standard deviation. We utilize widely accepted measures of clustering performance for our evaluation indexes, including the normalized mutual information (NMI) [40], the adjusted rand index (ARI) [41], and the clustering accuracy index (ACC) [42]. Both NMI and ACC range from 0 to 1, with 0 indicating the worst, and 1 representing the best. ARI has a range from -1 to 1, in which -1 means two data clusterings are completely dissimilar, 0 indicates that data clusterings are essentially random, and 1 means two data clusterings are perfectly aligned.
FKM, PFKM, csiFKM, siibFKM, and FWPKM employ a typical fuzzifier value of $m=2$ [43]. The fuzzifier value of MEFC is set to $\lambda=1$. For EKM, we set $\alpha=1$ for the four artificial datasets. For the 16 real-world datasets with varying feature numbers, the parameter $\alpha$ is set proportional to the data variance, as follows
$\alpha=2/\bar{d}_{0n},$ (46)
where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points. The reason for such selection is explained in Section IV-C.
V-B Artificial Datasets
Fig. 4 displays scatter plots of the four artificial datasets along with selected data clusterings. The average and the standard deviation of evaluation indexes over 50 trials are provided in Table II, where the best and the second-best results are highlighted in bold for easy comparison. On Data-A, the data points are sampled from Gaussian distributions. While on Data-B the data points are sampled from uniform distributions. Data-C and Data-D are a mixture of Gaussian and uniform distributions. The difference is that the majority of data points on Data-C are sampled from a Gaussian distribution while the majority of data points on Data-D are sampled from a uniform distribution. Additionally, Data-D has more classes than Data-C. All four artificial datasets are highly imbalanced with CV values greater than one.
We can see from Table II that the performance of benchmark algorithms is poor on these artificial datasets, especially on Data-C and Data-D. Nevertheless, the proposed EKM is effective, noticeably outperforming the best benchmark algorithm. The csiFKM has a certain effectiveness on Data-A and Data-B but does not exceed EKM. As seen in Fig. 4, the benchmark algorithms erroneously divide the majority class into multiple clusters to balance the data size (i.e., the uniform effect) while EKM does not. This experiment also proves that EKM is versatile in handling different data distributions.
Figure 4: Scatter diagrams of artificial datasets. (a)-(d) Primitive scatter diagrams with reference class labels (indicated by colors). (e)-(f) Clustering results of some benchmark algorithms. (i)-(l) Clustering results of the proposed EKM. The black crosses represent centroids obtained by each algorithm.
TABLE II: Experimental results on four artificial datasets with $CV>1$ (the best and second-best performances are in bold)
Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
Data-A NMI 0.5150 $\pm$0.0001 0.4982 $\pm$0.0001 0.2270 $\pm$0.0002 0.4984 $\pm$0.0001 0.7579 $\pm$0.0471 0.4924 $\pm$0.0503 0.1427 $\pm$0.0488 0.9126 $\pm$0.0000
ARI 0.2906 $\pm$0.0002 0.2485 $\pm$0.0002 0.0756 $\pm$0.0003 0.2489 $\pm$0.0003 0.8545 $\pm$0.0329 0.2974 $\pm$0.0490 0.0116 $\pm$0.0357 0.9616 $\pm$0.0000
ACC 0.6893 $\pm$0.0003 0.6075 $\pm$0.0006 0.4818 $\pm$0.0003 0.6089 $\pm$0.0009 0.9429 $\pm$0.0164 0.7443 $\pm$0.0192 0.4562 $\pm$0.0393 0.9933 $\pm$0.0000
Data-B NMI 0.5193 $\pm$0.0000 0.5160 $\pm$0.0000 0.2121 $\pm$0.0001 0.5181 $\pm$0.0001 0.7989 $\pm$0.0677 0.5090 $\pm$0.0419 0.1591 $\pm$0.0577 0.9981 $\pm$0.0057
ARI 0.2529 $\pm$0.0000 0.2452 $\pm$0.0000 0.0547 $\pm$0.0000 0.2498 $\pm$0.0003 0.8794 $\pm$0.0406 0.2461 $\pm$0.0486 0.0332 $\pm$0.0497 0.9992 $\pm$0.0023
ACC 0.6116 $\pm$0.0000 0.5852 $\pm$0.0002 0.4607 $\pm$0.0016 0.6022 $\pm$0.0009 0.9424 $\pm$0.0194 0.6304 $\pm$0.0278 0.4767 $\pm$0.0671 0.$9$$\pm$0.0004
Data-C NMI 0.0909 $\pm$0.0000 0.0797 $\pm$0.0000 0.0892 $\pm$0.0000 0.0782 $\pm$0.0001 0.1069 $\pm$0.0928 0.0769 $\pm$0.1281 0.0783 $\pm$0.0151 0.9514 $\pm$0.0000
ARI 0.0181 $\pm$0.0000 0.0034 $\pm$0.0001 0.0159 $\pm$0.0001 0.0015 $\pm$0.0001 0.0728 $\pm$0.1242 0.0199 $\pm$0.1713 0.0053 $\pm$0.0090 0.9807 $\pm$0.0000
ACC 0.5754 $\pm$0.0000 0.5302 $\pm$0.0002 0.5689 $\pm$0.0002 0.5237 $\pm$0.0002 0.6981 $\pm$0.2108 0.8207 $\pm$0.0436 0.5371 $\pm$0.0272 0.9987 $\pm$0.0000
Data-D NMI 0.4799 $\pm$0.0316 0.3185 $\pm$0.0010 0.3963 $\pm$0.0670 0.2353 $\pm$0.0554 0.3984 $\pm$0.0440 0.3022 $\pm$0.0729 0.1733 $\pm$0.0268 0.9463 $\pm$0.0104
ARI 0.1096 $\pm$0.0195 0.0467 $\pm$0.0007 0.1506 $\pm$0.0360 0.0389 $\pm$0.0183 0.0813 $\pm$0.0266 0.0339 $\pm$0.0260 0.0083 $\pm$0.0068 0.9954 $\pm$0.0049
ACC 0.3593 $\pm$0.0555 0.2559 $\pm$0.0080 0.4893 $\pm$0.0133 0.3699 $\pm$0.0380 0.3357 $\pm$0.0564 0.2330 $\pm$0.0398 0.2206 $\pm$0.0325 0.9763 $\pm$0.0045
V-C Study of Parameter Impact
We investigate the influence of the parameter $\alpha$ on EKM in this experiment. We perform EKM with $\alpha$ values of $0.1,0.2,0.5,0.8,1,2,5,8,10$ on the four artificial datasets. Fig. 5 presents the corresponding NMI values for these $\alpha$ values. The NMI value of HKM is considered as a reference. It is evident that when $\alpha>5$, the NMI of EKM closely aligns with that of HKM. This similarity arises because a sufficiently large $\alpha$ makes the objective of EKM nearly identical to the objective of HKM. As the value of $\alpha$ decreases, the NMI of EKM increases abruptly, which implies the emergence of repulsive force between centroids overcoming the uniform effect. When $\alpha<0.2$, EKM becomes inferior to HKM. This is because the centroids of EKM tend to overlap when $\alpha$ is particularly small (repulsion becomes attraction, see Fig. 3d when $\alpha=0.5$). Fortunately, the range of $\alpha$ that causes the centroids to overlap is narrow.
Figure 5: NMI of EKM with different $\alpha$ values on the four artificial datasets. The blue curve is the NMI of EKM, and the red line is the NMI of HKM as a reference.
V-D Real Datasets with Balanced Data
To evaluate EKM’s effectiveness on balanced data, we perform it on six real datasets: IS, Seeds, Heart Disease, Wine, Rice, and WDBC. These datasets are routinely used in the field of machine learning and are characterized by uniform class sizes, CV values less than 0.4, and are categorized as balanced datasets as per our protocol. Table III shows the clustering results. EKM delivers superior performance on the Wine dataset and is comparable to the baseline performance on the remaining five datasets. Although EKM’s mechanism makes it more advantageous for imbalanced data, this experiment proves that EKM is competitive for balanced data.
TABLE III: Experimental results on selected six real-world datasets with $CV<0.4$ (The best performance is in bold)
Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
IS NMI 0.5873 $\pm$0.0008 0.5950 $\pm$0.0057 0.6206 $\pm$0.0009 0.4962 $\pm$0.0241 0.5195 $\pm$0.0317 0.5925 $\pm$0.0371 0.5452 $\pm$0.0664 0.6618 $\pm$0.0138
ARI 0.4608 $\pm$0.0005 0.4960 $\pm$0.0059 0.4979 $\pm$0.0005 0.3650 $\pm$0.0149 0.3472 $\pm$0.0587 0.3776 $\pm$0.0794 0.4022 $\pm$0.0884 0.4810 $\pm$0.0089
ACC 0.5456 $\pm$0.0003 0.6578 $\pm$0.0146 0.5943 $\pm$0.0002 0.5276 $\pm$0.0139 0.4977 $\pm$0.0670 0.5840 $\pm$0.0511 0.5576 $\pm$0.0775 0.5609 $\pm$0.0118
Seeds NMI 0.7279 $\pm$0.0000 0.7318 $\pm$0.0056 0.7496 $\pm$0.0016 0.7170 $\pm$0.0000 0.7200 $\pm$0.0077 0.6893 $\pm$0.0015 0.5409 $\pm$0.0336 0.7315 $\pm$0.0000
ARI 0.7733 $\pm$0.0000 0.7768 $\pm$0.0058 0.7968 $\pm$0.0017 0.7607 $\pm$0.0000 0.7640 $\pm$0.0082 0.7143 $\pm$0.0034 0.4724 $\pm$0.0513 0.7715 $\pm$0.0000
ACC 0.9190 $\pm$0.0000 0.9209 $\pm$0.0023 0.9285 $\pm$0.0007 0.9143 $\pm$0.0000 0.9156 $\pm$0.0033 0.8952 $\pm$0.0015 0.7069 $\pm$0.0649 0.9190 $\pm$0.0000
Heart Disease NMI 0.3162 $\pm$0.0000 0.0142 $\pm$0.0132 0.2500 $\pm$0.0000 0.2326 $\pm$0.0975 0.1548 $\pm$0.0398 0.0364 $\pm$0.0476 0.0868 $\pm$0.0820 0.2571 $\pm$ 0.0000
ARI 0.3641 $\pm$0.0000 0.0033 $\pm$0.0042 0.3140 $\pm$0.0000 0.2995 $\pm$0.1230 0.0509 $\pm$0.0172 0.0163 $\pm$0.0523 0.1030 $\pm$0.1050 0.3162 $\pm$ 0.0000
ACC 0.8020 $\pm$0.0000 0.5249 $\pm$0.0202 0.7805 $\pm$0.0000 0.7653 $\pm$0.0689 0.6101 $\pm$0.0284 0.5354 $\pm$0.0549 0.6344 $\pm$0.0898 0.7815 $\pm$0.0000
Wine NMI 0.8759 $\pm$0.0000 0.8759 $\pm$0.0000 0.8759 $\pm$0.0000 0.8097 $\pm$0.0171 0.6966 $\pm$0.1663 0.5250 $\pm$0.0803 0.3976 $\pm$0.0794 0.8920 $\pm$0.0000
ARI 0.8975 $\pm$0.0000 0.8975 $\pm$0.0000 0.8975 $\pm$0.0000 0.8249 $\pm$0.0142 0.6283 $\pm$0.2549 0.3954 $\pm$0.1252 0.3518 $\pm$0.0987 0.9134 $\pm$0.0000
ACC 0.9663 $\pm$0.0000 0.9663 $\pm$0.0000 0.9663 $\pm$0.0000 0.9396 $\pm$0.0051 0.7797 $\pm$0.1834 0.7129 $\pm$0.0900 0.6476 $\pm$0.0742 0.9719 $\pm$0.0000
Rice NMI 0.5685 $\pm$0.0000 0.5688 $\pm$0.0000 0.5682 $\pm$0.0004 0.5742 $\pm$0.0008 0.5544 $\pm$0.0001 0.5437 $\pm$0.0032 0.0536 $\pm$0.0268 0.5659 $\pm$0.0005
ARI 0.6815 $\pm$0.0000 0.6824 $\pm$0.0000 0.6817 $\pm$0.0004 0.6876 $\pm$0.0008 0.6634 $\pm$0.0001 0.6471 $\pm$0.0109 0.0338 $\pm$0.0134 0.6777 $\pm$0.0004
ACC 0.9129 $\pm$0.0000 0.9131 $\pm$0.0000 0.9129 $\pm$0.0001 0.9147 $\pm$0.0002 0.9074 $\pm$0.0000 0.9023 $\pm$0.0033 0.5941 $\pm$0.0155 0.9117 $\pm$0.0001
WBDC NMI 0.5547 $\pm$0.0000 0.5612 $\pm$0.0000 0.5547 $\pm$0.0000 0.5700 $\pm$0.0000 0.3516 $\pm$0.2432 0.4202 $\pm$0.1246 0.1138 $\pm$0.0626 0.5513 $\pm$0.0025
ARI 0.6707 $\pm$0.0000 0.6829 $\pm$0.0000 0.6707 $\pm$0.0000 0.6895 $\pm$0.0000 0.4093 $\pm$0.3070 0.5250 $\pm$0.1734 0.0281 $\pm$0.0409 0.6444 $\pm$0.0028
ACC 0.9104 $\pm$0.0000 0.9139 $\pm$0.0000 0.9104 $\pm$0.0000 0.9156 $\pm$0.0000 0.7989 $\pm$0.1380 0.8557 $\pm$0.0810 0.5910 $\pm$0.0441 0.9027 $\pm$0.0009
V-E Real Datasets with Imbalanced Data
The considered datasets here are Ecoli, Htru2, Zoo, Glass, Shill Bidding, Anuran Calls, Occupancy Detection, Machine Failure, Pulsar Cleaned, and Bert-Embedded Spam. Among them, Htru2 and Occupancy Detection have a large amount of data, Glass and Ecoli have 6 and 8 classes respectively, and Zoo has high dimensionality. They have varying class sizes with CV values greater than 0.7. Except for Occupancy Detection and Zoo, the other eight datasets have CV values greater than 1. Bert-Embedded Spam has over 700 features. Since the Euclidean distance is ineffective for clustering algorithms on high-dimensional data, we execute principal component analysis on the Bert-Embedded Spam data. The first five principal components are used for clustering as adding more does not improve and even compromise performance. Table IV displays the results, where the best and the second-best results are in bold.
As can be seen from Table IV, in general, EKM obtains the best performances on eight datasets, and achieves the second-best clustering performances on the remaining two datasets (Zoo and Anuran Calls), as confirmed by all three measured clustering indexes. Specifically, although csiFKM outperforms EKM on the Zoo dataset, it performs substantially worse on Htru2, Occupancy Detection, Shill Bidding, etc. HKM performs well on Anuran Calls but underperforms on Shill Bidding, Machine Failure, and Bert-Embedded Spam. Due to the structural complexity of real datasets, it is reasonable that some algorithms perform well on certain datasets. In comparison, EKM performs consistently well on all datasets, especially on the Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam datasets, where it significantly outperforms the benchmark algorithms. Fig. 6 illustrates the scatter diagrams of Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam datasets, along with some clustering results. Comparing Fig. LABEL:sub@fig:htru2_fwpkm - LABEL:sub@fig:spam_mefc with Fig. LABEL:sub@fig:htru2_ekm - LABEL:sub@fig:spam_ekm, again, we can see the uniform effect exists in the benchmark algorithms. Conversely, the centroids obtained by EKM are farther away from each other, implying the existence of the repulsive force overcoming the uniform effect. Note that, we can only see one centroid in some figures because centroids overlap on the selected two coordinates. Overall, the experiment results provide strong evidence to support the superiority and consistent performance of EKM on imbalanced data. Note that, we simply set $\alpha$ proportional to the data variance. EKM’s performance can be further enhanced by fine-tuning the parameter $\alpha$.
TABLE IV: Experimental results on selected ten real-world datasets with $CV>0.7$ (the best and second-best performances are in bold)
Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
Zoo NMI 0.8381 $\pm$0.0268 0.7764 $\pm$0.0019 0.8179 $\pm$0.0000 0.6611 $\pm$0.0806 0.8657 $\pm$0.0054 0.7625 $\pm$0.0415 0.6652 $\pm$0.0804 0.7912 $\pm$0.0188
ARI 0.7546 $\pm$0.0487 0.6235 $\pm$0.0010 0.6444 $\pm$0.0000 0.4868 $\pm$0.0829 0.8715 $\pm$0.0034 0.6423 $\pm$0.1030 0.4743 $\pm$0.0948 0.8181 $\pm$0.0783
ACC 0.8248 $\pm$0.0310 0.6832 $\pm$0.0000 0.7228 $\pm$0.0000 0.6267 $\pm$0.0536 0.8703 $\pm$0.0078 0.7305 $\pm$0.0831 0.5958 $\pm$0.0604 0.8507 $\pm$0.0566
Glass NMI 0.3140 $\pm$0.0046 0.3073 $\pm$0.0003 0.3120 $\pm$0.0012 0.0721 $\pm$0.0069 0.3390 $\pm$0.0549 0.3424 $\pm$0.0413 0.1654 $\pm$0.0823 0.3764 $\pm$0.0370
ARI 0.1702 $\pm$0.0026 0.1529 $\pm$0.0005 0.1651 $\pm$0.0009 0.0070 $\pm$0.0059 0.1930 $\pm$0.0435 0.1648 $\pm$0.0366 0.0734 $\pm$0.0659 0.1978 $\pm$0.0193
ACC 0.4586 $\pm$0.0076 0.4021 $\pm$0.0011 0.4487 $\pm$0.0007 0.3368 $\pm$0.0054 0.4684 $\pm$0.0476 0.4202 $\pm$0.0270 0.4096 $\pm$0.0584 0.4796 $\pm$0.0086
Ecoli NMI 0.6379 $\pm$0.0040 0.5695 $\pm$0.0115 0.6096 $\pm$0.0076 0.5012 $\pm$0.0649 0.5985 $\pm$0.0241 0.5541 $\pm$0.0464 0.4746 $\pm$0.0424 0.6426 $\pm$0.0024
ARI 0.5032 $\pm$0.0070 0.4142 $\pm$0.0204 0.4815 $\pm$0.0270 0.3357 $\pm$0.0752 0.4807 $\pm$0.0898 0.3697 $\pm$0.0862 0.2832 $\pm$0.0500 0.5157 $\pm$0.0013
ACC 0.6461 $\pm$0.0103 0.5769 $\pm$0.0160 0.6238 $\pm$0.0218 0.5302 $\pm$0.0435 0.6276 $\pm$0.0711 0.5624 $\pm$0.0638 0.4598 $\pm$0.0410 0.6482 $\pm$0.0043
Htru2 NMI 0.4068 $\pm$0.0000 0.4100 $\pm$0.0004 0.4075 $\pm$0.0002 0.1289 $\pm$0.0002 0.0204 $\pm$0.0506 0.3671 $\pm$0.0290 0.0660 $\pm$0.0268 0.5872 $\pm$0.0003
ARI 0.6071 $\pm$0.0000 0.5796 $\pm$0.0005 0.6075 $\pm$0.0002 0.0462 $\pm$0.0002 0.0077 $\pm$0.0649 0.5704 $\pm$0.0379 -0.0117 $\pm$0.0074 0.7333 $\pm$0.0002
ACC 0.9366 $\pm$0.0000 0.9252 $\pm$0.0001 0.9366 $\pm$0.0000 0.6091 $\pm$0.0002 0.$8$$\pm$0.0637 0.9337 $\pm$0.0126 0.5170 $\pm$0.0151 0.9661 $\pm$0.0000
Shill Bidding NMI 0.0021 $\pm$0.0001 0.0023 $\pm$0.0000 0.0042 $\pm$0.0002 0.0012 $\pm$0.0000 0.0295 $\pm$0.0359 0.0351 $\pm$0.0047 0.0105 $\pm$0.0330 0.6216 $\pm$0.0000
ARI -0.0006 $\pm$0.$0$0.0002 $\pm$0.$0$0.0002 $\pm$0.$0$0.0002 $\pm$0.$0$0.0485 $\pm$0.0482 -0.0854 $\pm$0.0118 0.0079 $\pm$0.0252 0.7914 $\pm$0.0000
ACC 0.5020 $\pm$0.0002 0.5059 $\pm$0.0000 0.5090 $\pm$0.0003 0.5035 $\pm$0.0002 0.7457 $\pm$0.0961 0.7869 $\pm$0.0182 0.5937 $\pm$0.0750 0.9674 $\pm$0.0000
Anuran Calls NMI 0.6775 $\pm$0.0195 0.5699 $\pm$0.0002 0.6155 $\pm$0.0022 0.4270 $\pm$0.0153 0.5575 $\pm$0.0516 0.5178 $\pm$0.0339 0.5159 $\pm$0.0247 0.6229 $\pm$0.0134
ARI 0.5826 $\pm$0.0174 0.3414 $\pm$0.0009 0.4017 $\pm$0.0027 0.2254 $\pm$0.0694 0.4329 $\pm$0.1363 0.2368 $\pm$0.0694 0.4268 $\pm$0.0500 0.5145 $\pm$0.0329
ACC 0.6441 $\pm$0.0134 0.4363 $\pm$0.0009 0.5103 $\pm$0.0073 0.3705 $\pm$0.0230 0.5682 $\pm$0.0799 0.4101 $\pm$0.0348 0.5123 $\pm$0.0271 0.5821 $\pm$0.0420
Occupancy Detection NMI 0.4890 $\pm$0.0001 0.4720 $\pm$0.0003 0.4710 $\pm$0.0002 0.0539 $\pm$0.0007 0.0934 $\pm$0.1560 0.2982 $\pm$0.1379 0.0591 $\pm$0.0464 0.5389 $\pm$0.0002
ARI 0.5959 $\pm$0.0001 0.5699 $\pm$0.0004 0.5670 $\pm$0.0002 0.0209 $\pm$0.0002 0.0521 $\pm$0.2248 0.2175 $\pm$0.2226 0.0363 $\pm$0.0682 0.6750 $\pm$0.0002
ACC 0.8912 $\pm$0.0000 0.8825 $\pm$0.0001 0.8815 $\pm$0.0001 0.5772 $\pm$0.0004 0.7597 $\pm$0.0596 0.7026 $\pm$0.1361 0.6117 $\pm$0.0595 0.9163 $\pm$0.0001
Machine Failure NMI 0.0248 $\pm$0.0000 0.0281 $\pm$0.0000 0.0310 $\pm$0.0000 0.0284 $\pm$0.0003 0.2285 $\pm$0.0569 0.0540 $\pm$0.0292 0.0137 $\pm$0.0161 0.5825 $\pm$0.1869
ARI -0.0116 $\pm$0.$0$0.0083 $\pm$0.$0$0.0052 $\pm$0.0000 0.0007 $\pm$0.0001 0.2217 $\pm$0.0510 -0.0037 $\pm$0.0382 0.0031 $\pm$0.0083 0.6506 $\pm$0.1987
ACC 0.5673 $\pm$0.0000 0.5418 $\pm$0.0002 0.5202 $\pm$0.0001 0.5146 $\pm$0.0007 0.9653 $\pm$0.0543 0.6125 $\pm$0.0412 0.5481 $\pm$0.0358 0.9863 $\pm$0.0068
Pulsar Cleaned NMI 0.0311 $\pm$0.0002 0.0167 $\pm$0.0000 0.0201 $\pm$0.0000 0.0236 $\pm$0.0011 0.0523 $\pm$0.0813 0.0044 $\pm$0.0081 0.0090 $\pm$0.0067 0.0544 $\pm$0.0001
ARI 0.0370 $\pm$0.0003 0.0069 $\pm$0.0000 0.0131 $\pm$0.0001 -0.0011 $\pm$0.0002 0.0731 $\pm$0.1193 -0.0067 $\pm$0.0160 -0.0039 $\pm$0.0080 0.1082 $\pm$0.0002
ACC 0.7329 $\pm$0.0008 0.5714 $\pm$0.0001 0.6171 $\pm$0.0004 0.5019 $\pm$0.0012 0.9488 $\pm$0.0973 0.9043 $\pm$0.1231 0.6036 $\pm$0.0659 0.8834 $\pm$0.0002
Bert-Embedded Spam NMI 0.1517 $\pm$0.0001 0.1249 $\pm$0.0000 0.0472 $\pm$0.0374 0.1122 $\pm$0.0023 0.0330 $\pm$0.0791 0.0431 $\pm$0.0946 0.0238 $\pm$0.0004 0.6704 $\pm$0.0000
ARI 0.0830 $\pm$0.0003 0.0550 $\pm$0.0000 0.0289 $\pm$0.0265 -0.0026 $\pm$0.0007 0.0080 $\pm$0.1084 0.0296 $\pm$0.1287 -0.0669 $\pm$0.0004 0.8216 $\pm$0.0000
ACC 0.6453 $\pm$0.0002 0.6174 $\pm$0.0000 0.5837 $\pm$0.0390 0.5482 $\pm$0.0012 0.8210 $\pm$0.0816 0.8281 $\pm$0.0832 0.8038 $\pm$0.0002 0.9664 $\pm$0.0000
Figure 6: Scatter diagrams of some selected real imbalanced datasets (from left to right: Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam). The first two features are used as coordinates for visualization. (a)-(d) Primitive scatter diagrams with reference class labels. (e)-(f) Clustering results of some benchmark algorithms. (i)-(l) Clustering results of the proposed EKM. The black crosses represent centroids obtained by each algorithm.
V-F Empirical Convergence and Algorithm Efficiency
Table V shows the average number of iterations and the average time per run (each algorithm is run 5000 times). The numbers in brackets in the table are the ranking numbers of the corresponding algorithms. AVK and MDK give the average and median number of iterations and computational time of each algorithm on all test datasets, respectively.
As evident from Table V, AVK and MDK are consistent except that EKM is faster than FKM in AVK and slower in MDK. This is because FKM is slow on the Bert-Embedded Spam dataset. HKM is the fastest algorithm as it does not require membership calculation. MEFC comes second, followed by FKM and EKM. EKM is 30% faster than FKM in AVK and slower than FKM by 9% in MDK. In terms of absolute speed, MEFC, FKM, and EKM operate within the same order of magnitude. This aligns with their identical time complexity of $O(NK^{2}P)$ per iteration, where $N$ is the instance number, $K$ is the cluster number, and $P$ is the data dimension or feature number. The table also demonstrates the empirical convergence of EKM, averaging 44 iterations or a median of 20 iterations to reach the specified convergence (45). In summary, EKM has a low computational cost.
TABLE V: Average number of iterations and average calculation time of each algorithm
Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
Data-A Iter 17.3 13.7 29.0 33.5 30.7 28.0 44.6 12.1
Time 0.005(1) 0.008(3) 0.013(4) 0.033(7) 0.021(6) 0.277(8) 0.016(5) 0.007(2)
Data-B Iter 9.2 13.2 30.1 28.8 15.1 17.4 24.1 17.5
Time 0.003(1) 0.008(2) 0.013(6) 0.031(7) 0.011(5) 0.168(8) 0.010(4) 0.009(3)
Data-C Iter 16.3 17.6 36.0 23.9 35.0 42.1 73.7 23.5
Time 0.008(1) 0.014(2) 0.023(4) 0.048(7) 0.040(6) 0.624(8) 0.035(5) 0.017(3)
Data-D Iter 14.8 23.6 25.4 39.8 24.5 25.2 7.3 15.6
Time 0.033(2) 0.091(5) 0.082(4) 0.364(7) 0.142(6) 1.867(8) 0.030(1) 0.058(3)
IS Iter 15.9 40.1 20.9 53.0 49.8 32.1 29.6 24.1
Time 0.021(1) 0.110(5) 0.058(2) 0.218(7) 0.104(4) 0.824(8) 0.152(6) 0.074(3)
Seeds Iter 8.5 10.4 11.1 13.3 12.6 26.0 32.5 12.0
Time 0.001(1) 0.001(3) 0.001(2) 0.003(6) 0.002(4) 0.028(8) 0.005(7) 0.002(5)
Heart Disease Iter 9.9 269.3 12.5 22.7 500.0 498.8 3.3 18.7
Time 0.002(1) 0.074(5) 0.003(3) 0.081(6) 2.095(8) 1.646(7) 0.002(2) 0.005(4)
Wine Iter 7.0 13.5 10.6 25.3 15.7 40.4 19.4 11.7
Time 0.000(1) 0.002(4) 0.001(2) 0.004(6) 0.002(5) 0.037(8) 0.005(7) 0.001(3)
Rice Iter 8.4 8.7 8.9 12.7 14.0 17.8 72.8 9.1
Time 0.004(1) 0.009(4) 0.006(2) 0.021(6) 0.011(5) 0.214(8) 0.072(7) 0.009(3)
WDBC Iter 7.4 10.8 10.6 32.4 500.0 500.0 16.2 9.9
Time 0.001(1) 0.003(4) 0.002(2) 0.009(5) 2.307(8) 0.933(7) 0.013(6) 0.002(3)
Zoo Iter 4.4 31.0 23.8 35.0 34.4 37.8 2.3 498.9
Time 0.001(1) 0.005(4) 0.003(3) 0.009(6) 0.006(5) 0.048(7) 0.001(2) 0.069(8)
[…]